please help me solve this equation
x+y=-3
x-y=1

Answer by CharlesG2(834)

please help me solve this equation
x+y=-3
x-y=1
x + y = -3 --> x = -3 - y (subsitute this into 2nd equation)
-3 - y - y = 1
-3 - 2y = 1
-2y = 4
y = -2
x + (-2) = -3
x = -3 + 2
x = -1